# Consistency of variational inference for Besov priors in non-linear inverse problems

### v1

@@ 5.1. Relations between information distances and the $L^{2}$ norm
$D_{\rho}(P_{1}\|P_{2})=\left\{\begin{aligned} &\frac{1}{\rho-1}\log\int\bigg{(}{\frac{dP_{1}}{dP_{2}}}\bigg{)}^{\rho-1}dP_{1}&\text{if}\ P_{1}\ll P_{2},\\ &+\infty&\text{otherwise}.\end{aligned}\right$.
@@
Moreover, the Rényi divergence $D_{\rho}(P_{1}||P_{2})$ is a non-decreasing function of $\rho$, which particularly gives $D(P_{1}\|P_{2})\leq D_{2}(P_{1}\|P_{2}).$ The Hellinger distance $h$ is defined as
@@ 5.2. Contraction rates for rescaled Besov priors
We are going to verify the three condition in [49, Theorem 2.1]. Steps (i) to (iii) below verify conditions (C1) to (C3) directly.
@@
$H_{N}(\varepsilon)=\Big{\{}{\theta:\theta=\theta_{1}+\theta_{2},\lVert{\theta_{1}}\rVert_{(H^{k})^{*}}\leq M^{l}\varepsilon_{N}/L_{\mathcal{G}}(Mr_{N}(\varepsilon)),\|\theta_{2}\|_{B_{pp}^{\alpha}}\leq M(\varepsilon/\varepsilon_{N})^{\frac{2}{p}}}\Big{\}}$
@@
(i) For (C1), we follow the method used in the proof of [15, Theorem 7.1.4]. Let
@@
The definition of $H_{N}(\varepsilon)$ implies that a $\frac{\bar{m}\varepsilon}{C_{L}(Mr_{N}(\varepsilon))^{l}}$-covering in $\lVert{\ \cdot\ }\rVert_{H^{-\kappa}(\mathcal{Z})}$ of $B_{pp}^{\alpha}(M(\varepsilon/\varepsilon_{N})^{\frac{2}{p}})$ is a $\frac{2\bar{m}\varepsilon}{C_{L}(Mr_{N}(\varepsilon))^{l}}$-covering in $\lVert{\ \cdot\ }\rVert_{H^{-\kappa}(\mathcal{Z})}$ of $\Theta_{N}(\varepsilon)$ for $\bar{m}\geq M^{l}$. It is sufficient to consider
@@
Choose a minimum finite set $S_{j}^{{}^{\prime}}$ of points in each set $S_{j}$ such that every $\theta\in S_{j}$ is within Hellinger distance $j\bar{m}\varepsilon$ of at least one of these points. By metric entropy bound above, for $j$ fixed, there are at most $N(j\varepsilon)$ such points $\theta_{jl}\in S_{j}^{{}^{\prime}}$, and from [15, Corollary 7.1.3] for each $\theta_{jl}$ there exists a test $\Psi_{N,jl}$ such that
@@
$h(p_{\theta},p_{\theta_{0}})$ $\geq\frac{1}{2C_{U}\cdot(M\varepsilon/\varepsilon_{N})^{p}}\|\mathcal{G}(\theta)-\mathcal{G}(\theta_{0})\|_{L^{2}_{\lambda}(\mathcal{X},V)}$
$\geq\frac{1}{2C_{U}C_{T}\cdot(M\varepsilon/\varepsilon_{N})^{p+q}}F(\|f_{\theta}-f_{\theta_{0}}\|).$
From the inequality and direct calculations, we note that
$\{\theta\in\Theta_{N}(\varepsilon):h(p_{\theta},p_{\theta_{0}})\geq 4\bar{m}\varepsilon\}\supseteq\{\theta\in\Theta_{N}(\varepsilon):N\varepsilon_{N}^{\frac{2p}{p+1}}\|\mathcal{G}(\theta)-\mathcal{G}(\theta_{0})\|_{L^{2}_{\lambda}}^{\frac{2}{p+1}}\geq\tilde{C}_{1}N\varepsilon^{2}\}$
@@
$\{\theta\in\Theta_{N}(\varepsilon):h(p_{\theta},p_{\theta_{0}})\geq 4\bar{m}\varepsilon\}\supseteq\{\theta\in\Theta_{N}(\varepsilon):N\varepsilon_{N}^{\frac{2p+2q}{p+q+1}}[F(\|f_{\theta}-f_{\theta_{0}}\|)]^{\frac{2}{p+q+1}}\geq C_{1}N\varepsilon^{2}\}$
for $\tilde{C}_{1}=(8\bar{m}C_{U}M^{p})^{\frac{2}{p+1}}$ and $C_{1}=(8\bar{m}C_{U}C_{T}M^{p+q})^{\frac{2}{p+q+1}}$. Combined with (5.7), we have
$\mathop{\sup}_{\scriptsize\begin{array}[]{c}{\theta\in\Theta_{N}(\varepsilon)}\\ {L(P^{(N)}_{\theta},P^{(N)}_{\theta_{0}})\geq N\varepsilon^{2}}\end{array}}P^{(N)}_{\theta}(1-\Psi_{N})\leq\exp\{-CN\varepsilon^{2}\}$ (5.8)
@@
$L(P^{(N)}_{\theta},P^{(N)}_{\theta_{0}}):=N\varepsilon_{N}^{\frac{2p+2q}{p+q+1}}[F(\|f_{\theta}-f_{\theta_{0}}\|)]^{\frac{2}{p+q+1}}/C_{1}\text{ or }N\varepsilon_{N}^{\frac{2p}{p+1}}\|\mathcal{G}(\theta)-\mathcal{G}(\theta_{0})\|_{L^{2}_{\lambda}}^{\frac{2}{p+1}}/\tilde{C}_{1}.$
@@
$\Pi_{N}(\|\theta\|_{\mathcal{R}}>M(\varepsilon/\varepsilon_{N})^{\frac{2}{p}})\leq\Pi^{\prime}(\|\theta^{\prime}\|_{\mathcal{R}^{\prime}}>cM(N\varepsilon^{2})^{\frac{1}{p}})\leq c_{1}\exp(-c_{2}M^{p}N\varepsilon^{2})\leq\frac{1}{2}\exp(-CN\varepsilon^{2})$
@@
$\geq\exp(-\frac{1}{2}N\varepsilon_{N}^{2}\|\theta_{0}\|_{B_{pp}^{\alpha}}^{p})\cdot\Pi^{\prime}(\theta:\lVert{\theta}\rVert_{(H^{\kappa})^{*}}\leq C_{\mathcal{G}}({\bar{M}})\varepsilon_{N}(N\varepsilon_{N}^{2})^{\frac{1}{p}},\|\theta\|_{\mathcal{R}^{\prime}}\leq M^{\prime}(N\varepsilon_{N}^{2})^{\frac{1}{p}})$
$\geq\exp(-\frac{1}{2}N\varepsilon_{N}^{2}\|\theta_{0}\|_{B_{pp}^{\alpha}}^{p})\cdot\Big{(}\Pi^{\prime}(\lVert{\theta}\rVert_{(H^{\kappa})^{*}}\leq C_{\mathcal{G}}({\bar{M}})\varepsilon_{N}(N\varepsilon_{N}^{2})^{\frac{1}{p}})-\Pi^{\prime}(\|\theta\|_{\mathcal{R}^{\prime}}>M^{\prime}(N\varepsilon_{N}^{2})^{\frac{1}{p}})\Big{)}$
@@
$P_{\theta_{0}}^{(N)}\hat{Q}N\varepsilon_{N}^{\frac{2p}{p+1}}\|\mathcal{G}(\theta)-\mathcal{G}(\theta_{0})\|_{L^{2}_{\lambda}}^{\frac{2}{p+1}}\lesssim N(\varepsilon_{N}^{2}+\gamma^{2}_{N}),$
@@
$P_{\theta_{0}}^{(N)}\hat{Q}N\varepsilon_{N}^{\frac{2p+2q}{p+q+1}}[F(\|f_{\theta}-f_{\theta_{0}}\|)]^{\frac{2}{p+q+1}}\lesssim N(\varepsilon_{N}^{2}+\gamma^{2}_{N}).$
@@
$P_{\theta_{0}}^{(N)}\hat{Q}\|\mathcal{G}(\theta)-\mathcal{G}(\theta_{0})\|_{L^{2}_{\lambda}}^{\frac{2}{p+1}}\lesssim\varepsilon_{N}^{\frac{2}{p+1}}+\gamma^{2}_{N}\cdot\varepsilon_{N}^{-\frac{2p}{p+1}},$
@@
$P_{\theta_{0}}^{(N)}\hat{Q}[F(\|f_{\theta}-f_{\theta_{0}}\|)]^{\frac{2}{p+q+1}}\lesssim\varepsilon_{N}^{\frac{2}{p+q+1}}+\gamma^{2}_{N}\cdot\varepsilon_{N}^{-\frac{2p+2q}{p+q+1}}.$
@@
The probability measure $Q_{N}$ is defined as the push-forward of $\tilde{Q}_{N}$ via $\Psi$, that is, $Q_{N}=\tilde{Q}_{N}\circ\Psi^{-1}.$ It is easy to see $Q_{N}\in{\mathcal{Q}}_{G}$. We first consider the upper bound of $\frac{1}{N}D(Q_{N}\|\Pi_{N})$. Because the KL divergence decreases under push-forward [45, section 10], we have
@@
Then, it is sufficient to consider the upper bound of $D(Exp(p;\theta_{0,lr},\tau_{l})\|Exp(p;0,\sigma_{l}))$. When $l\leq J$, we have
@@
For the last term, by the wavelet characterization of Sobolev norms (see [15, section 4]), we deduce that
@@
$\lesssim$ $E\big{[}\big{(}\lVert{\theta_{0}}\rVert^{2l}_{B_{\gamma\gamma}^{\beta^{\prime}}(\mathcal{Z})}+\lVert{\sum_{l=-1}^{+\infty}\sum_{r\in R_{l}}\tau_{l}Z_{lr}\chi\psi_{lr}}\rVert^{2l}_{B_{\gamma\gamma}^{\beta^{\prime}}(\mathcal{Z})}\big{)}$
@@
where $\theta_{lr}=0$ for all $l>J$. Using (1.3) and wavelet characterization of Sobolev norms (see [15, section 4]), we deduce that
@@ 5.3. Contraction rates for high-dimensional Besov priors
Following the proof of Theorem 3.3, we will now verify conditions (C1) through (C3) as outlined in [49, Theorem 2.1].
@@
$\log N\Big{(}B_{H^{\tilde{\alpha}}}\big{(}{M\sqrt{N}\varepsilon}/{(\sqrt{N}\varepsilon_{N})^{\frac{2}{p}}}\big{)},\|\cdot\|_{H^{-\kappa}(\mathcal{Z})},\frac{\bar{m}\varepsilon}{C_{L}(Mr_{N}(\varepsilon))^{l}}\leq\exp\Big{\{}{-cN\varepsilon^{2}}\Big{\}}\Big{)}$
@@
$\log N(B_{H^{\tilde{\alpha}}}\big{(}{M\sqrt{N}\varepsilon}/{(\sqrt{N}\varepsilon_{N})^{\frac{2}{p}}}\big{)},\frac{\bar{m}\varepsilon}{C_{L}(Mr_{N}(\varepsilon))^{l}})$
@@
$\leq$ $C_{E}\left(\frac{M^{l+1}C_{L}}{\bar{m}}\right)^{d/(\tilde{\alpha}+{\kappa})}\cdot(\frac{\varepsilon}{\varepsilon_{N}})^{\frac{2d(l+1)}{p(\tilde{\alpha}+\kappa)}}\cdot\Big{(}\frac{(\sqrt{N}\varepsilon)^{1-\frac{2}{p}}}{\varepsilon}\Big{)}^{\frac{d}{p(\tilde{\alpha}+\kappa)}}$
$\leq$ $C_{E}\cdot(\frac{\varepsilon}{\varepsilon_{N}})^{2}\cdot\big{(}N^{\frac{p-2}{2p}}\varepsilon_{N}^{-\frac{2}{p}}\big{)}^{\frac{d}{p(\tilde{\alpha}+\kappa)}}$
$=$ $C_{E}\cdot(\frac{\varepsilon}{\varepsilon_{N}})^{2}\cdot\big{(}(N\varepsilon_{N}^{2})^{\frac{2\alpha+2\kappa+d}{d}\cdot\frac{p-2}{2p}}\cdot(N\varepsilon_{N}^{2})^{\frac{\alpha+\kappa}{d}\cdot\frac{2}{p}}\big{)}^{\frac{d}{p(\tilde{\alpha}+\kappa)}}$
@@
$\geq$ $\Pi_{J}^{\prime}\bigg{(}\theta=\theta_{1}+\theta_{2}+\theta_{3}:\lVert{\theta_{1}}\rVert_{(H^{\kappa}({\mathbb{R}^{d}}))^{*}}\leq\frac{cM^{l}\varepsilon_{N}(N\varepsilon_{N}^{2})^{\frac{1}{p}}}{L_{\mathcal{G}}(Mr_{N}(\varepsilon))},\lVert{\theta_{2}}\rVert_{B_{pp}^{\alpha}({\mathbb{R}^{d}})}\leq\frac{c^{\prime}}{2}(MN\varepsilon^{2})^{\frac{1}{p}},$
@@
$\geq\Pi_{N}(\theta:\lVert{\mathcal{G}(\theta)-\mathcal{G}(P_{J}(\theta_{0}))}\rVert_{L^{2}_{\lambda}}\leq(C_{3}\exp\{-2U_{\mathcal{G}}^{2}(\bar{M})\}-C_{0})\varepsilon_{N},\lVert{\theta-\theta_{0}}\rVert_{\mathcal{R}}\leq M^{\prime})$
@@
The proof of this theorem is similar to that of Theorem 3.4, so we will omit same procedure. The notations $\tilde{\mathcal{Q}}_{E}^{J},\Psi_{J},{\mathcal{Q}}_{E}^{J}$ are defined identically as in (3.17-3.19). We recall that
@@
It is necessary to bound $\frac{1}{N}D(Q\|\Pi_{N})$ and $Q[D(P_{\theta_{0}}\|P_{\theta})]$ respectively. We define $\tilde{Q}_{N}$ to be $\mathop{\bigotimes}_{l=-1}^{J}\mathop{\bigotimes}_{r\in R_{l}}Exp(q;\theta_{0,lr},\tau),\quad\tau\simeq 2^{-J(\alpha+{\kappa}+d/2)},\quad\theta_{0,lr}=\langle{\theta_{0}},{\psi_{lr}}\rangle_{L^{2}(\mathcal{Z})}.$ The probability measure $Q_{N}$ is defined as the push-forward of $\tilde{Q}_{N}$ via $\Psi_{J}$, that is, $Q_{N}=\tilde{Q}_{N}\circ\Psi_{J}^{-1}.$ It is easy to see $Q_{N}\in{\mathcal{Q}}_{E}^{J}(q)$. We also see that the prior $\Pi_{N}$ can be represented as $\Pi_{N}=\tilde{\Pi}_{N}\circ\Psi_{J}^{-1},$ where $\tilde{\Pi}_{N}=\mathop{\bigotimes}_{l=-1}^{J}\mathop{\bigotimes}_{r\in R_{l}}Exp(p;0,\sigma_{l})$ for $\sigma_{l}=2^{-l(\alpha+\frac{d}{2}-\frac{d}{p})}(N\varepsilon_{N}^{2})^{-\frac{1}{p}}$. We first consider the upper bound of $\frac{1}{N}D(Q_{N}\|\Pi_{N})$. Because the KL divergence decreases under push-forward [45, section 10], we have
@@
$Q_{N}D(P_{\theta_{0}}\|P_{\theta})\lesssim$ $E\big{[}\big{(}\lVert{\theta_{0}}\rVert^{2l}_{B_{\gamma\gamma}^{\beta^{\prime}}(\mathcal{Z})}+\lVert{\sum_{l=-1}^{J}\sum_{r\in R_{l}}\tau Z_{lr}\chi\psi_{lr}}\rVert^{2l}_{B_{\gamma\gamma}^{\beta^{\prime}}(\mathcal{Z})}\big{)}$ (5.15)

### v2

@@ 6.1. Relations between information distances and the $L^{2}$ norm
$D_{\gamma}(P_{1}\|P_{2})=\left\{\begin{aligned} &\frac{1}{\rho-1}\log\int\bigg({\frac{dP_{1}}{dP_{2}}}\bigg)^{\rho-1}dP_{1}&\text{if}\ P_{1}\ll P_{2},\\ &+\infty&\text{otherwise}.\end{aligned}\right$.
@@
Moreover, the Rényi divergence $D_{\gamma}(P_{1}||P_{2})$ is a non-decreasing function of $\gamma$, which particularly gives $D(P_{1}\|P_{2})\leq D_{2}(P_{1}\|P_{2}).$ The Hellinger distance $h$ is defined as
@@ 6.2. Contraction rates for rescaled Besov priors
To deduce the contraction result, We verify the three conditions formulated in [56, Theorem 2.1]. Specifically, it is sufficient to prove that the following three conditions hold for a loss function $L(\cdot,\cdot)$ and constants $C,C_{1},C_{2},C_{3}>0$ with $C>C_{2}+C_{3}+2$:
  • (C1)
    For any $\varepsilon>\varepsilon_{N}$,there exists a set $\Theta_{N}(\varepsilon)$ and a testing function $\Psi_{n}$, such that
    $P^{(N)}_{\theta_{0}}\Psi_{N}+\sup_{\begin{subarray}{c}\theta\in\Theta_{N}(\varepsilon)\\ L(P^{(N)}_{\theta},P^{(N)}_{\theta_{0}})\geq N\varepsilon^{2}\end{subarray}}P^{(N)}_{\theta}(1-\Psi_{N})\leq\exp\{-CN\varepsilon^{2}\}$
  • (C2)
    For any $\varepsilon>\varepsilon_{N}$, the set $\Theta_{N}(\varepsilon)$ in (C1) satisfies $\Pi_{N}(\Theta_{N}(\varepsilon)^{c})\leq e^{-CN\varepsilon^{2}}$;
  • (C3)
    For some constant $\gamma>1$, $\Pi_{N}(\theta:D_{\gamma}(P^{(N)}_{\theta_{0}}\|P^{(N)}_{\theta})\leq C_{3}N\varepsilon_{N}^{2})\geq e^{-C_{2}N\varepsilon^{2}}$.
@@
We are going to verify the three condition in [56, Theorem 2.1]. Steps (i) to (iii) below verify conditions (C1) to (C3) directly.
@@
$H_{N}(\varepsilon)=\Big\{{\theta=\theta_{1}+\theta_{2}:\lVert{\theta_{1}}\rVert_{(H^{k})^{*}}\leq M^{l}\varepsilon_{N}/L_{\mathcal{G}}(Mr_{N}(\varepsilon)),\|\theta_{2}\|_{B_{pp}^{\alpha}}\leq M(\varepsilon/\varepsilon_{N})^{\frac{2}{p}}}\Big\}$
@@
(i) For (C1), we follow the method used in the proof of [18, Theorem 7.1.4]. Let
@@
The definition of $H_{N}(\varepsilon)$ implies that, for $\bar{m}\geq M^{l}$, a $\frac{\bar{m}\varepsilon}{C_{L}(Mr_{N}(\varepsilon))^{l}}$-covering in $\lVert{\ \cdot\ }\rVert_{H^{-\kappa}(\mathcal{Z})}$ of $B_{pp}^{\alpha}(M(\varepsilon/\varepsilon_{N})^{\frac{2}{p}})$ is a $\frac{2\bar{m}\varepsilon}{C_{L}(Mr_{N}(\varepsilon))^{l}}$-covering in $\lVert{\ \cdot\ }\rVert_{H^{-\kappa}(\mathcal{Z})}$ of $\Theta_{N}(\varepsilon)$. It is sufficient to consider
@@
Choose a minimum finite set $S_{j}^{{}^{\prime}}$ of points in each set $S_{j}$ such that every $\theta\in S_{j}$ is within Hellinger distance $j\bar{m}\varepsilon$ of at least one of these points. By metric entropy bound above, for $j$ fixed, there are at most $N(j\varepsilon)$ such points $\theta_{jl}\in S_{j}^{{}^{\prime}}$, and from [18, Corollary 7.1.3] for each $\theta_{jl}$ there exists a test $\Psi_{N,jl}$ such that
@@
$h(p_{\theta},p_{\theta_{0}})$ $\geq\frac{1}{2C_{U}\cdot(M\varepsilon/\varepsilon_{N})^{\mu}}\|\mathcal{G}(\theta)-\mathcal{G}(\theta_{0})\|_{L^{2}_{\lambda}(\mathcal{X},V)}$
$\geq\frac{1}{2C_{U}C_{T}\cdot(M\varepsilon/\varepsilon_{N})^{\mu+\nu}}F(\|f_{\theta}-f_{\theta_{0}}\|).$
From the inequality and direct calculations, we note that the set
$\{\theta\in\Theta_{N}(\varepsilon):h(p_{\theta},p_{\theta_{0}})\geq 4\bar{m}\varepsilon\}$
contains
$\{\theta\in\Theta_{N}(\varepsilon):N\varepsilon_{N}^{\frac{2\mu}{\mu+1}}\|\mathcal{G}(\theta)-\mathcal{G}(\theta_{0})\|_{L^{2}_{\lambda}}^{\frac{2}{\mu+1}}\geq\tilde{C}_{1}N\varepsilon^{2}\}$
@@
$\{\theta\in\Theta_{N}(\varepsilon):N\varepsilon_{N}^{\frac{2\mu+2\nu}{\mu+\nu+1}}[F(\|f_{\theta}-f_{\theta_{0}}\|)]^{\frac{2}{\mu+\nu+1}}\geq C_{1}N\varepsilon^{2}\}$
for $\tilde{C}_{1}=(8\bar{m}C_{U}M^{\mu})^{\frac{2}{\mu+1}}$ and $C_{1}=(8\bar{m}C_{U}C_{T}M^{\mu+\nu})^{\frac{2}{\mu+\nu+1}}$. Combined with (6.7), we have
$\sup_{\begin{subarray}{c}\theta\in\Theta_{N}(\varepsilon)\\ L(P^{(N)}_{\theta},P^{(N)}_{\theta_{0}})\geq N\varepsilon^{2}\end{subarray}}P^{(N)}_{\theta}(1-\Psi_{N})\leq\exp\{-CN\varepsilon^{2}\}$ (6.8)
@@
$L(P^{(N)}_{\theta},P^{(N)}_{\theta_{0}}):=N\varepsilon_{N}^{\frac{2\mu+2\nu}{v+1}}[F(\|f_{\theta}-f_{\theta_{0}}\|)]^{\frac{2}{\mu+\nu+1}}/C_{1}\text{ or }N\varepsilon_{N}^{\frac{2p}{\mu+1}}\|\mathcal{G}(\theta)-\mathcal{G}(\theta_{0})\|_{L^{2}_{\lambda}}^{\frac{2}{\mu+1}}/\tilde{C}_{1}.$
@@
$\Pi_{N}(\|\theta\|_{\mathcal{R}}>M(\varepsilon/\varepsilon_{N})^{\frac{2}{p}})\leq\Pi^{\prime}(\|\theta^{\prime}\|_{\mathcal{R}^{\prime}}>cM(N\varepsilon^{2})^{\frac{1}{p}})\leq c_{1}e^{-c_{2}M^{p}N\varepsilon^{2}}\leq\frac{1}{2}e^{-CN\varepsilon^{2}}$
@@
$\geq e^{-\frac{1}{p}N\varepsilon_{N}^{2}\|\theta_{0}\|_{B_{pp}^{\alpha}}^{p}}\cdot\Pi^{\prime}(\theta:\lVert{\theta}\rVert_{(H^{\kappa})^{*}}\leq C_{\mathcal{G}}({\bar{M}})\varepsilon_{N}(N\varepsilon_{N}^{2})^{\frac{1}{p}},\|\theta\|_{\mathcal{R}^{\prime}}\leq M^{\prime}(N\varepsilon_{N}^{2})^{\frac{1}{p}})$
$\geq e^{-\frac{1}{p}N\varepsilon_{N}^{2}\|\theta_{0}\|_{B_{pp}^{\alpha}}^{p}}\Big(\Pi^{\prime}(\lVert{\theta}\rVert_{(H^{\kappa})^{*}}\leq C_{\mathcal{G}}({\bar{M}})\varepsilon_{N}(N\varepsilon_{N}^{2})^{\frac{1}{p}})-\Pi^{\prime}(\|\theta\|_{\mathcal{R}^{\prime}}>M^{\prime}(N\varepsilon_{N}^{2})^{\frac{1}{p}})\Big)$
@@
$P_{\theta_{0}}^{(N)}\hat{Q}N\varepsilon_{N}^{\frac{2\mu}{\mu+1}}\|\mathcal{G}(\theta)-\mathcal{G}(\theta_{0})\|_{L^{2}_{\lambda}}^{\frac{2}{\mu+1}}\lesssim N(\varepsilon_{N}^{2}+\gamma^{2}_{N}),$
@@
$P_{\theta_{0}}^{(N)}\hat{Q}N\varepsilon_{N}^{\frac{2\mu+2\nu}{\mu+\nu+1}}[F(\|f_{\theta}-f_{\theta_{0}}\|)]^{\frac{2}{\mu+\nu+1}}\lesssim N(\varepsilon_{N}^{2}+\gamma^{2}_{N}).$
@@
$P_{\theta_{0}}^{(N)}\hat{Q}\|\mathcal{G}(\theta)-\mathcal{G}(\theta_{0})\|_{L^{2}_{\lambda}}^{\frac{2}{\mu+1}}\lesssim\varepsilon_{N}^{\frac{2}{\mu+1}}+\gamma^{2}_{N}\cdot\varepsilon_{N}^{-\frac{2\mu}{\mu+1}},$
@@
$P_{\theta_{0}}^{(N)}\hat{Q}[F(\|f_{\theta}-f_{\theta_{0}}\|)]^{\frac{2}{\mu+\nu+1}}\lesssim\varepsilon_{N}^{\frac{2}{\mu+\nu+1}}+\gamma^{2}_{N}\cdot\varepsilon_{N}^{-\frac{2\mu+2\nu}{\mu+\nu+1}}.$
@@
The probability measure $Q_{N}$ is defined as the push-forward of $\tilde{Q}_{N}$ via $\Psi$, that is, $Q_{N}=\tilde{Q}_{N}\circ\Psi^{-1}.$ It is easy to see $Q_{N}\in{\mathcal{Q}}_{E}$. We first consider the upper bound of $\frac{1}{N}D(Q_{N}\|\Pi_{N})$. Because the KL divergence decreases under push-forward [52, section 10], we have
@@
Then, it is sufficient to deduce the upper bound of $D(Exp(p;\theta_{0,lr},\tau_{l})\|Exp(p;0,\sigma_{l}))$. We note that $C_{p,b}$ denotes the normalization constant of the p-exponential distribution $Exp(p;0,b)$ (see Lemma 7.1 for the explicit expression). When $l\leq J$, we have
@@
For the last term, by the wavelet characterization of Sobolev norms (see [18, section 4]), we deduce that
@@
[…]
@@
where $\theta_{lr}=0$ for all $l>J$. Using (1.3) and wavelet characterization of Sobolev norms (see [18, section 4]), we deduce that
@@ 6.3. Contraction rates for high-dimensional Besov priors
Following the proof of Theorem 3.3, we will now verify conditions (C1) through (C3) as outlined in [56, Theorem 2.1].
@@
$\log N\Big(B_{H^{\tilde{\alpha}}}\big({M\sqrt{N}\varepsilon}/{(\sqrt{N}\varepsilon_{N})^{\frac{2}{p}}}\big),\|\cdot\|_{H^{-\kappa}(\mathcal{Z})},\frac{\bar{m}\varepsilon}{C_{L}(Mr_{N}(\varepsilon))^{l}}\Big)\leq\exp\Big\{{-cN\varepsilon^{2}}\Big\}$
@@
$\log N(B_{H^{\tilde{\alpha}}}\big({M\sqrt{N}\varepsilon}/{(\sqrt{N}\varepsilon_{N})^{\frac{2}{p}}}\big),{\color[rgb]{0,0,0}\|\cdot\|_{H^{-\kappa}(\mathcal{Z})},}\frac{\bar{m}\varepsilon}{C_{L}(Mr_{N}(\varepsilon))^{l}})$
@@
$\leq$ $C_{E}\left(\frac{M^{l+1}C_{L}}{\bar{m}}\right)^{d/(\tilde{\alpha}+{\kappa})}\cdot(\frac{\varepsilon}{\varepsilon_{N}})^{\frac{2d(l+1)}{p(\tilde{\alpha}+\kappa)}}\cdot\Big(\frac{(\sqrt{N}\varepsilon)^{1-\frac{2}{p}}}{\varepsilon}\Big)^{\color[rgb]{0,0,0}\frac{d}{\tilde{\alpha}+\kappa}}$
$\leq$ $C_{E}\cdot(\frac{\varepsilon}{\varepsilon_{N}})^{2}\cdot\big(N^{\frac{p-2}{2p}}\varepsilon_{N}^{-\frac{2}{p}}\big)^{\color[rgb]{0,0,0}\frac{d}{\tilde{\alpha}+\kappa}}$
$=$ $C_{E}\cdot(\frac{\varepsilon}{\varepsilon_{N}})^{2}\cdot\big((N\varepsilon_{N}^{2})^{\frac{2\alpha+2\kappa+d}{d}\cdot\frac{p-2}{2p}}\cdot(N\varepsilon_{N}^{2})^{\frac{\alpha+\kappa}{d}\cdot\frac{2}{p}}\big)^{\color[rgb]{0,0,0}\frac{d}{\tilde{\alpha}+\kappa}}$
@@
$\geq$ $\Pi_{J}^{\prime}\bigg(\theta_{1}+\theta_{2}+\theta_{3}:\lVert{\theta_{1}}\rVert_{(H^{\kappa}({\mathbb{R}^{d}}))^{*}}\leq\frac{cM^{l}\varepsilon_{N}(N\varepsilon_{N}^{2})^{\frac{1}{p}}}{L_{\mathcal{G}}(Mr_{N}(\varepsilon))},\lVert{\theta_{2}}\rVert_{B_{pp}^{\alpha}({\mathbb{R}^{d}})}\leq\frac{c^{\prime}}{2}(MN\varepsilon^{2})^{\frac{1}{p}},$
$\qquad\qquad\qquad\qquad\quad\lVert{\theta_{3}}\rVert_{H^{\alpha+\frac{d}{2}-\frac{d}{p}}({\mathbb{R}^{d}})}\leq\frac{c}{2}\sqrt{M}\sqrt{N}\varepsilon,\theta_{i}\in\mathrm{span}\{\psi_{lr}\},i=1,2,3\bigg).$
@@
$\geq\Pi_{N}(\theta:\lVert{\mathcal{G}(\theta)-\mathcal{G}(P_{J}(\theta_{0}))}\rVert_{L^{2}_{\lambda}}\leq C^{\prime}(\bar{M})\varepsilon_{N},\lVert{\theta-P_{J}(\theta_{0})}\rVert_{\mathcal{R}}\leq M^{\prime\prime})$
@@
The proof of this theorem is similar to that of Theorem 3.4, so we will omit same procedure. The notations $\tilde{\mathcal{Q}}_{E}^{J},\Psi_{J},{\mathcal{Q}}_{E}^{J}$ are defined identically as in (3.16-3.18). We recall that
@@
It is necessary to bound $\frac{1}{N}D(Q\|\Pi_{N})$ and $Q[D(P_{\theta_{0}}\|P_{\theta})]$ respectively. We define $\tilde{Q}_{N}$ to be $\mathop{\bigotimes}_{l=-1}^{J}\mathop{\bigotimes}_{r\in R_{l}}Exp(q;\theta_{0,lr},\tau),\quad\tau\simeq 2^{-J(\alpha+{\kappa}+d/2)},\quad\theta_{0,lr}=\langle{\theta_{0}},{\psi_{lr}}\rangle_{L^{2}(\mathcal{Z})}.$ The probability measure $Q_{N}$ is defined as the push-forward of $\tilde{Q}_{N}$ via $\Psi_{J}$, that is, $Q_{N}=\tilde{Q}_{N}\circ\Psi_{J}^{-1}.$ It is easy to see $Q_{N}\in{\mathcal{Q}}_{E}^{J}(q)$. We also see that the prior $\Pi_{N}$ can be represented as $\Pi_{N}=\tilde{\Pi}_{N}\circ\Psi_{J}^{-1},$ where $\tilde{\Pi}_{N}=\mathop{\bigotimes}_{l=-1}^{J}\mathop{\bigotimes}_{r\in R_{l}}Exp(p;0,\sigma_{l})$ for $\sigma_{l}=2^{-l(\alpha+\frac{d}{2}-\frac{d}{p})}(N\varepsilon_{N}^{2})^{-\frac{1}{p}}$. We first consider the upper bound of $\frac{1}{N}D(Q_{N}\|\Pi_{N})$. Because the KL divergence decreases under push-forward [52, section 10], we have
@@
[…]